# Detection and Measurement of Syntactic Templates in Generated Text

### v1

@@ 8 Style Memorization
Past work has shown that models sometimes memorize portions of pre-training data. Carlini et al. (2022) show a lower bound of 1% verbatim memorized data. We next show how our syntactic template analysis can be used to evaluate how much models memorize from pre-training data, beyond strict token sequence matches (RQ 3).
@@ 8.2 Results
We randomly sample 10k documents and look at the fraction of memorized outputs based on exact-match and the POS sequence (“style”) memorization. We average the fraction memorized over 1,000 seeds for sampling the datasets.
@@
The implications of our looser definition of memorization allows us to capture instances of memorization where exact tokens may be substituted during generation, but where an output span is nonetheless structurally the same as a source string. In Figure 7, we provide examples of substitutions that occur that are not captured by exact-memorization definitions, yet demonstrate that the particular style of that training point has been memorized. We find that these cases often include synonym swaps, or different numbers being generated.
@@ 9 Conclusions
In this work, we introduce syntactic templates as a framework for analyzing subtle repetitive characteristics in model-generated text. We show that this analysis can also extend to human-written references and downstream tasks, and find that the pre-training data contains many of these identified templates. We show that this definition of repetition is useful for detecting subtle types of data “memorization”. Our hope is that this work inspires additional research into characterizing where (in data) observed stylistic patterns in LLM outputs originate.
@@
Second, we use third party tools to tag our text abstractions; however these tools are not deterministic and can contain errors in the tags they assign to sequences, particularly if a sequence contains text from another language. We assume that the majority of the text we analyze is in English, and that any errors are superseded by the frequency of common templates.
@@ Acknowledgements
This work was supported in part by the National Science Foundation (NSF) grants IIS 2211954, IIS 2145479, IIS 2107524.

### v2

@@ 8 Style Memorization
Past work has shown that models memorize portions of pre-training data. Carlini et al. (2022) show a lower bound of 1% verbatim memorized data. We next show how our syntactic template analysis can be used to evaluate how much models memorize from pre-training data, beyond strict token sequence matches (RQ3).
@@ 8.2 Results
We randomly sample 10k documents and look at the fraction of memorized outputs based on exact-match and the POS sequence (“style”) memorization using the diversity package (e.g., Fig. 2). We average the fraction memorized over 1,000 seeds for sampling the datasets.
@@
The implications of our looser definition of memorization allows us to capture instances of memorization where exact tokens may be substituted during generation, but where an output span is nonetheless structurally the same as a source string. Note that this method will by default also capture duplicate text in addition to softly memorized sequence. In Figure 8, we provide sampled examples of substitutions that occur that are not captured by exact-memorization definitions, yet demonstrate that the particular style of that training point has been memorized. We find that these cases often include synonym swaps, or different numbers being generated.
@@ 9 Conclusions
In this work, we introduce syntactic templates as a framework for analyzing subtle repetitive characteristics in model-generated text. We show that this analysis can also extend to human-written references and downstream tasks, and find that the pre-training data contains many of these identified templates. We show that evaluating repetition in parts-of-speech sequences is useful for detecting subtle types of data “memorization”. Our hope is that this work inspires additional research into characterizing where (in data) observed stylistic patterns in LLM outputs originate.
@@
Second, we use third party tools to tag our text abstractions; however these tools are deterministic, but can contain errors in the tags they assign to sequences, particularly if a sequence contains text from another language. We assume that the majority of the text we analyze is in English, and that any errors are superseded by the frequency of common templates.
@@ Acknowledgements
We thank Niloofar Mireshghallah for guidance on the memorization experiments. We also thank Kyle Lo, and Luca Soldaini for their advice and assistance on accessing OLMo and pretraining data. This work was supported in part by the National Science Foundation (NSF) grants IIS 2211954, IIS 2145479, IIS 2107524.